OpenStudy (anonymous):

What is the period, in seconds, of a simple pendulum of length 2 meters? Use the gravitational constant g = 9.8 m/s^2 and round your answer to two decimal places.

OpenStudy (anonymous):

2.84 is what I have?

OpenStudy (anonymous):

\[T = 2\pi \sqrt{\frac{L}{G}}\] Substitute the values and get T

OpenStudy (anonymous):

T=2π∗ √L/g T=2π∗ √2/9.8 T=2π∗ (0.4517539515) T=2.84s